OpenStudy (anonymous):

The formula for the amount with simple interest is A = P(1 + rt). Find r, if A = 224, P = 200, and t = 2.

OpenStudy (anonymous):

Solve for r: Divide by P A/P=1+rt Subtract 1 \[\frac{A}{P}-1=rt\] Divide by t. \[\frac{(\frac{A}{P}-1)}{t}=r\] Then just plug in the numbers and punch it in a calculator :P